Equations are given whose graphs enclose a region. Find the area of the region. (Give an exact answer. Do not round.)

y =

1
2

x2;  y = x2 − 2x
Answer #1

fined the area of region We need to enclosed by J = the curves y = 12x² and 2 = 80 - 22 2 Y=lax 2 = x-2x 48 121 Panteration fU 2= -2 and o 11 2 Y = 12x² 12xo o and 12x * ( 2 ) ie y = 0 and 48 121 A Note : fa g@ X x=a x = 6 Area enclosed between two c: ا ه (as) - (21)) - 22 - la C 11 (-11&² - 2x) dx از 0 3 2 ع . ] - اند و 3 2 0 3 2. 2 ) - (()-( رال 1 0 2 8 (23| ا lx + [2] HThank you.